# A Comprehensive Reanalysis of K2-18 b’s JWST NIRISS+NIRSpec Transmission Spectrum

### v1

@@ 3 NIRSpec Data Reductions
The K2-18 system was also observed by JWST’s Near Infrared Spectrograph (NIRSpec; Ferruit et al., 2012; Birkmann et al., 2014) instrument using the G395H grating from 18:37:38 UTC on January 20, 2023 to 01:11:32 UTC on January 21, 2023 as part of the same JWST GO Program as the NIRISS data.This observation, which also covered one full transit of K2-18 b, used the Bright Object Time Series (BOTS) mode with the F290LP filter, SUB2048 subarray, and NRSRAPID readout pattern. The spectra were dispersed over two detectors: NRS1, whose wavelength range covers 2.73–3.72 m, and NRS2, whose wavelength range covers 3.82–5.17 m. In total, the NIRSpec observation covers 2.73–5.17 m with a gap from 3.72–3.82 m.
We perform a total of four NIRSpec reductions: one with the FIREFLy pipeline, one with the ExoTEDRF pipeline, and two with the Eureka! pipeline. We summarize the most relevant choices made in each reduction, as well as a comparison to the reduction choices made in Madhusudhan et al. (2023)’s JExoRES NIRSpec reduction, in Table 2. As with NIRISS, we show several additional diagnostic figures for our reductions in Appendix C.
@@
| Trim | First 575 integrations in | None | None | First 25 integrations | First 5 minutes |
@@
| Error bar inflation | Red noise via Allan | Jitter term | White noise term | White noise term | ? |
| variance | |||||
@@
Note. — “Default” refers to the default handling of the step in the jwst pipeline. The system parameters (, , and ) are the same for each reduction, and are calculated in Madhusudhan et al. (2023). This is not an exhaustive list of reduction steps. ‘?’ indicates reduction configuration settings not mentioned in Madhusudhan et al. (2023).
@@ 3.1 FIREFLy
We also used the FIREFLy pipeline to conduct a full time-series analysis of the K2-18 NIRSpec/G395H data. Starting with the uncal data files, we use Stage 1 of the jwst pipeline for group-level detector and instrument corrections. We apply the data quality initialization and saturation steps to the uncal data files, but like the FIREFLy reduction of May & MacDonald et al. (2023) (and unlike the FIREFLy reduction of Moran & Stevenson et al. 2023) we do not apply a superbias scaling step. Instead, we use the default jwst superbias file. We also use the default background subtraction. We apply the reference pixel correction and linearity step, but not the dark current step. We also skip the jump step to avoid cosmic ray false positives due to a less well-constrained ramp linear slope, as typically occurs with data sets with fewer than 25 groups per integration. We apply a group-level 1/f subtraction and finish Stage 1 with the standard ramp fitting and gain-scale steps. Moving to Stage 2 of the jwst pipeline, we next apply integration-level detector and instrument corrections. We skip the flat fielding step, only applying the assign World Coordinate System (WCS) step before moving to Stage 3, the FIREFLy pipeline’s stellar extraction step.
We clean bad pixels, which we determine by flagging pixels with sharp variance spikes of over 100 and other known bad pixels in NIRSpec G395H, using lacosmic (van Dokkum, 2001) as the first part of the stellar extraction. We apply an integration-level noise subtraction with a PSF mask, measure x- and y-shifts, and extract the 1D stellar spectrum as a box extraction using a standard NIRSpec/G395H trace. We use an aperture full-width of 5.2 pixels for NRS1 and 2.41 pixels for NRS2, as they minimize the scatter of the out-of-transit white light curve.
@@
We also use batman to fit the cleaned spectrophotometric light curves. We bin the data in four ways: resolving powers , , and , and at native pixel resolution. For each binning regime, we trim the first 575 columns of NRS1, first 8 columns of NRS2, and last 18 columns of NRS2. We fit for the mid-transit time T0, scaled semimajor axis a/R∗, and impact parameter b in NRS1 and NRS2, and find that the weighted average orbital parameters , , and are consistent with those reported by Madhusudhan et al. (2023). We therefore fix the orbital parameters to theirs and fit the white light curve for the radius ratio Rp/R∗, quadratic limb darkening parameters u+ and u-, and a linear term in time using the emcee Python package. For consistency with Madhusudhan et al. (2023), we experiment with including both a linear and quadratic term in time, but we find that only a linear term is favored; we therefore choose to only include a linear term. We show the results of our white light curve fit in Figure 2.
@@
As with NIRISS, we use the lmfit Python package to perform the Levenberg-Marquardt algorithm to fit the radius ratio for the spectrophotometric light curves. For this fit we fix the limb darkening parameters and linear systematic term to the fitted white light curve values in addition to the already fixed orbital parameters. We elect to fix the spectrophotometric limb darkening values to the white light curve values as limb darkening is not expected to be strongly wavelength dependent at the redder wavelengths of NIRSpec (May & MacDonald et al. 2023; Moran & Stevenson et al. 2023). We account for red noise by inflating error bars in quadrature for wavelength bins where the Allan variance is above log linear. For most points this does not result in error bar inflation, and for points that are inflated, it is typically by a factor of no more than two. We show the spectroscopic light curve fits and residuals for each bin of our reduction in Appendix B. With both NIRISS and NIRSpec data in hand, we compare the low-resolution data-variants from each reduction code in Figure 3 ( for NIRISS; for NIRSpec) and directly compare our pixel-level data, binned to the same resolution as Madhusudhan et al. (2023), in Figure 4.
@@
To reduce the NIRSpec/G395H observations, we also use the exoTEDRF pipeline (Radica et al., 2023; Feinstein et al., 2023; Radica, 2024), which has recently been updated to support NIRSpec observations (Radica, 2024, Ahrer et al. submitted). We start from the raw, uncalibrated data files and follow the procedure outlined in Ahrer et al. (submitted). For completeness, we summarize the key points here. We perform standard saturation and superbias corrections on the data frames (e.g., Alderson et al., 2023) before correcting the background and 1/ noise at the group-level, using the median method and a trace mask width of 16 pixels. As with SOSS, we perform a time-domain outlier flagging (Radica et al., 2024b) with a rejection threshold of 12. We then repeat the background and 1/ correction at the integration-level, after ramp fitting, in order to remove any remaining traces of the background flux. We then interpolate all pixels which have a non-zero data quality flag, or are flagged as 10 temporal outliers using a median of the surrounding pixels in space and time, respectively. For the spectral extraction, we use the edgetrigger algorithm (Radica et al., 2022b) to determine the centroids of the NRS1 and NRS2 spectral traces, and then perform a box aperture extraction with a width of eight pixels.
We follow the same light curve fitting procedure as with the SOSS light curves — jointly fitting the NRS1 and NRS2 white light curves, sharing the planet’s orbital parameters between the two and independently fitting for chromatic parameters like the scaled planet radius, limb darkening and systematics. There is no spot crossing in the NIRSpec light curves, and so we do not include a Gaussian spot model. However, we keep the linear slope and error inflation term added in quadrature to the flux errors that were included in the SOSS fits. In all, our white light curve fits have 15 free parameters. We again use emcee for the fits, using the same amount of chains, steps, and burn in as above.
As with SOSS, we fix the orbital parameters to the values used by Madhusudhan et al. (2023) (there is, once again, no difference in the resulting spectra using the Radica et al. (2022a) orbital solution), and fit the spectrophotometric light curves at four different resolutions, the pixel level (that is, one light curve per detector column), , , and .
@@ 3.3 Eureka! Reductions
For our first Eureka! reduction, which we henceforth refer to as reduction “A”, we start with the uncalibrated files and run Stage 1 and Stage 2 which are wrapped around the jwst pipeline. We use the default steps in both stages with the exception of the jump step and the photom step, which we skip. We also use Eureka!’s group-level background subtraction and a custom bias scale factor (using a smoothing filter with a window length of 30 and group 1) in Stage 1444See also Moran & Stevenson et al. 2023.
@@
In Stage 3 we extract the stellar time-series spectra. First we correct the curved trace of the spectra and perform a background subtraction where we used the median of the entire frame, excluding the area within 5 pixels of the center of the trace. Prior to the calculation of the median background, we perform two iterations of outlier masking using thresholds of 3 along the time and 5 times the median along the spatial axis. We use an aperture of 7 pixels, corresponding to a 3-pixel half width, to extract the time-series spectra with the optimal extraction (Horne, 1986).
@@
We generate pixel-level and binned light curves in Stage 4 of Eureka! following the same binning scheme as FIREFLy. We use a 3-clipping at a 10-pixel-rolling-median to reduce the number of outliers. We fit the light curves by fixing the system parameters to the ones reported in Madhusudhan et al. (2023) to allow for direct comparison and the MCMC python package emcee. We opt to use a quadratic limb-darkening law, where we fix one parameter () to limb parkening parameters generated using the exotic-ld Python package (Grant & Wakeford, 2024) and the MPS-ATLAS (set 2) stellar atmosphere model grid (Kostogryz et al., 2023). We use a simple linear trend in time to account for a systematic trend in our light curves.
@@
Our second Eureka! reduction, which we will henceforth refer to as reduction “B”, used version 0.11.dev446+gf5d684ee.d20240712; we started with the _uncal files produced with SDP_VER 2023_3b, and we used version 1.15.1 of the jwst pipeline with CRDS version 11.17.26 and CRDS context jwst_1252. Our reduction procedure largely followed the Eureka! reduction described in Schlawin et al. (2024); we summarize the important details below.
@@
We increase the stage 1 jump step rejection threshold to 15, use Eureka!’s custom bias correction (computed using a smoothing filter of length 15 on group 1), and subtract the median background flux per column and per group after having first masked pixels near the spectral trace. In stage 2, we skip the flat fielding, photometric calibration, and extract_1d steps. In stage 3, we first mask all pixels labeled ‘DO_NOT_USE’ in the data quality array, and we then straighten the trace so that the center of the stellar PSF is centered on the subarray (using only integer-pixel shifts). We then 4 clip background pixels along the time axis and 2.5 along the spatial axis and then subtract the mean background flux per column and per integration computed using only the pixels 6 px away from the center of the spectral trace. Next, we perform optimal spectral extraction (Horne, 1986) using a cleaned median integration to compute the spatial profile with a 4-pixel half width. In stage 4, we manually mask several pixel columns (18 in NRS1 and 9 in NRS2) that exhibit unusually high noise levels compared to their neighbors, likely due to unmasked bad pixels in an earlier stage. We then spectrally bin the data into multiple different spectral resolutions and mask any integration that was a 4 outlier with respect to a 20-integration wide boxcar filter, likely due to a missed cosmic ray.
@@
When fitting the light curves, our astrophysical model consists of a batman transit model with quadratic limb-darkening coefficients fixed to a PHOENIX limb-darkening model computed by exotic-ld (Husser et al., 2013). We first verify that our broadband light curves are consistent with the orbital parameters reported by Madhusudhan et al. (2023), and then we fix our orbital parameters to those of Madhusudhan et al. (2023). Our systematic model consists of a linear trend in time and a linear correlation with the position and the width of the spatial PSF. We also fit for a white noise multiplier to account for any excess white noise in the data beyond the estimated photon noise. We trimmed the first 25 integrations to remove any initial detector settling. To estimate the best fit model values and their corresponding uncertainties, we use the dynesty dynamic nested sampling algorithm (Speagle, 2020) with 121 live points, ‘multi’ bounds, the ‘rwalk’ sampling algorithm, and a convergence criterion of , where is the Bayesian evidence.
@@ 4.1 Retrieval Configuration
We employ two open source retrieval codes, POSEIDON and BeAR, to explore the range of atmospheric properties consistent with the revised K2-18 b’s transmission spectra derived above. We additionally show retrieval results using the data from Madhusudhan et al. (2023) in Appendix A. Our retrieval configurations are constructed to encompass the parameter space explored by Madhusudhan et al. (2023), while also considering additional factors (such as planetary mass uncertainties and including the 2nd order NIRISS SOSS data). Our model settings, retrieval configurations, and a comparison with Madhusudhan et al. (2023) are summarized in Table 3.
@@
| (citation) | Benneke et al. (2019) | Benneke et al. (2017) | Benneke et al. (2019) |
@@ 4.1.1 POSEIDON
We first retrieved K2-18 b’s transmission spectrum using the retrieval code POSEIDON555https://github.com/MartianColonist/POSEIDON (MacDonald & Madhusudhan, 2017; MacDonald, 2023). POSEIDON is widely applied to interpret JWST transmission spectra, including hot Jupiters (e.g., Grant et al., 2023; Fournier-Tondreau et al., 2024b), sub-Neptunes (Piaulet-Ghorayeb et al., 2024), super-Earths (e.g., Moran & Stevenson et al. 2023; May & MacDonald et al. 2023), and terrestrial worlds (e.g., Lim et al., 2023; Cadieux et al., 2024). Here we use POSEIDON to conduct a systematic retrieval survey of the sensitivity of K2-18 b’s atmospheric inferences to 60 different data variants (reduction codes and/or data spectral resolution). Our POSEIDON retrieval configuration is constructed to be as close as possible to that described in Madhusudhan et al. (2023), allowing a fair comparison, but we emphasize and justify below several areas where we adopted different model choices.
@@
Our atmosphere model for K2-18 b assumes a H2-He background gas (with a solar ratio of He/H2 = 0.17; Asplund et al. 2009), commensurate with its planetary radius and mass ( and , respectively). We construct atmospheres ranging from –10 bar with 100 layers spaced uniformly in log-pressure under hydrostatic equilibrium. We adopt bar as the lowest atmosphere pressure (vs. the bar limit implied by Madhusudhan et al. 2023) to reduce the risk of strong bands saturating the atmosphere. We solve hydrostatic equilibrium using a boundary condition where the freely retrieved reference pressure corresponds to the radial distance where (the radius from Benneke et al. 2019). The atmospheric pressure-temperature (P-T) profile is freely retrieved following the prescription in Madhusudhan & Seager (2009). We consider the same potential molecular species as Madhusudhan et al. (2023), which includes the major O-, C-, and N-bearing species expected at K2-18 b’s equilibrium temperature (255 K; Benneke et al. 2019): H2O, CH4, NH3, HCN, CO, and CO2; as well as several disequilibrium species proposed as biosignatures in cool sub-Neptunes (Madhusudhan et al., 2021): C2H6S (DMS), CS2, CH3Cl, OCS, and N2O. We fit the volume mixing ratios of these gases independently, with their number densities then determined via the ideal gas law. Given the large uncertainties on K2-18 b’s mass (; Cloutier et al. 2019), we prescribe the planet’s mass as a free parameter using a Gaussian prior (Madhusudhan et al. 2023 fixed K2-18 b’s mass), allowing mass uncertainties to propagate into the retrieved atmospheric composition (Batalha et al., 2019a). Finally, we consider aerosols following the inhomogeneous cloud and haze parameterization of MacDonald & Madhusudhan (2017).
@@
We calculate model spectra by solving the equation of radiative transfer in a cylindrical coordinate system for 100 incident stellar rays attenuated by atmospheric opacity along the line of sight. The POSEIDON opacity database recently underwent a major update in POSEIDON v1.2 (Mullens et al., 2024), including a state-of-the-art new ExoMol CH4 line list (Yurchenko et al., 2024). Our POSEIDON opacities are pre-computed at a high resolution ( = 0.01 cm-1, equivalent to at 1 ) for a grid of temperatures and pressures using the open source Python package Cthulhu666https://github.com/MartianColonist/Cthulhu (Agrawal & MacDonald, 2024). Our K2-18 b retrievals use opacities derived from the following line lists/measured cross sections: H2O (POKAZATEL; Polyansky et al. 2018), CH4 (MM; Yurchenko et al. 2024), NH3 (CoYuTe; Coles et al. 2019), HCN (Harris; Barber et al. 2014, CO (Li2015; Li et al. 2015), CO2 (UCL-4000; Yurchenko et al. 2020), C2H6S (DMS) (HITRAN measured xsec; Sharpe et al. 2004) CS2 (HITRAN-2020; Gordon et al. 2022), CH3Cl (HITRAN-2020; Gordon et al. 2022), OCS (OYT8; Owens et al. 2024), and N2O (HITEMP; Hargreaves et al. 2019). We further include continuum collision-induced absorption from H2-H2 and H2-He pairs (Karman et al., 2019) and Rayleigh scattering for all gases (MacDonald & Lewis, 2022). We calculate model transmission spectra using opacity sampling on a wavelength grid from 0.58–5.3 at a spectral resolution of 100,000. We adopt this exceptionally high spectral resolution for all our POSEIDON retrievals for three reasons: (i) to ensure negligible opacity sampling errors when interpreting pixel-level data; (ii) to consistently sample cross sections for every data combination (even at lower resolutions); and (iii) to reduce the risk of opacity accuracy differences between our work and Madhusudhan et al. (2023) (since the latter did not state their assumed sampling resolution). While our 100,000 models impose a greater computational burden than is strictly necessary for retrievals of JWST data777Through the JWST Transiting Exoplanet Early Release Science Program, we found a useful heuristic that 20,000 suffices for = 100 and 60,000 suffices for 3,000 (pixel-level)., this ensures high-accuracy in our retrieved atmospheric properties.
@@
Our POSEIDON retrievals additionally account for stellar contamination from unocculted active regions, commonly known as the transit light source effect (Rackham et al., 2018). We employ a one-heterogeneity model (e.g., Rathcke et al., 2021; Lim et al., 2023; Fournier-Tondreau et al., 2024a), defined by the stellar photosphere temperature, the heterogeneity temperature (corresponding to spots/faculae if the heterogeneity is cooler/warmer than the photosphere), and the heterogeneity covering fraction. The stellar contamination contribution is calculated by interpolating PHOENIX models (Husser et al., 2013) using the PyMSG package (Townsend & Lopez, 2023). For fixed stellar properties, we adopt the stellar radius from Benneke et al. (2019) () and the stellar surface gravity, metallicity, and effective temperature from Tabernero et al. (2024) ([Fe/H] = , = 4.93, = 3563 K). These updated K2-18 stellar properties were published after Madhusudhan et al. (2023), so we adopt them for a more accurate characterization of the transit light source effect.

### v2

@@ 3 NIRSpec Data Reductions
The K2-18 system was also observed by JWST’s Near Infrared Spectrograph (NIRSpec; Ferruit et al., 2012; Birkmann et al., 2014) instrument using the G395H grating as part of JWST GO Program 2722 (PI: N. Madhusudhan - Madhusudhan, 2023). As with NIRISS, the details of these observations are discussed in Madhusudhan et al. (2023). We perform a total of four NIRSpec reductions: one with the FIREFLy pipeline, one with the ExoTEDRF pipeline, and two with the Eureka! pipeline. We bin the data in four ways: resolving powers , , and , and at native pixel resolution. We summarize the most relevant choices made in each reduction, as well as a comparison to the reduction choices made in Madhusudhan et al. (2023)’s JExoRES NIRSpec reduction, in Table 2. Similarly with NIRISS, we show the point-by point discrepancies between reductions in Appendix C. Combining these reductions with our NIRISS reductions, we compare the low-resolution data-variants from each reduction code in Figure 3 ( for NIRISS; for NIRSpec) and directly compare our pixel-level data, binned to the same resolution as Madhusudhan et al. (2023), in Figure 4.
@@
| Trim | First 575 pixels in | None | None | First 25 integrations | First 5 minutes |
@@
| Error bar inflation | Red noise via residual | Jitter term | White noise term | White noise term | ? |
| RMS-bin size trend | |||||
@@
Note. — “Default” refers to the default handling of the step in the jwst pipeline. The system parameters (, , and ) are the same for each reduction, and are calculated in Madhusudhan et al. (2023). The listed orbital parameters were used to either confirm consistency with Madhusudhan et al. (2023) (FIREFLy, Eureka! B) or to show negligible differences between resulting spectra with an alternative orbital solution (exoTEDRF, from Radica et al., 2022a). Note that this is not an exhaustive list of reduction steps. ‘?’ indicates reduction configuration settings not mentioned in Madhusudhan et al. (2023).
@@ 3.1 FIREFLy
We also used the FIREFLy pipeline to conduct a full time-series analysis of the K2-18 NIRSpec/G395H data. We again list a more thorough inventory of the steps in Table 2, describing here the most pertinent choices made. We skip the jump step in Stage 1 to avoid cosmic ray false positives due to a less well-constrained ramp linear slope, as typically occurs with data sets with fewer than 25 groups per integration. In Stage 3, we clean bad pixels by flagging ones with sharp variance spikes of over 100 using lacosmic (van Dokkum, 2001) and other known bad pixels in NIRSpec G395H as the first part of the stellar extraction. Our box extraction’s aperture full-widths were selected as they minimize the scatter of the out-of-transit white light curve.
@@
In the light curve fitting step, we trim the first 575 pixel columns of NRS1, first 8 pixel columns of NRS2, and last 18 pixel columns of NRS2. We find that our fitted weighted average orbital parameters are consistent with those reported by Madhusudhan et al. (2023); we therefore fix the orbital parameters to theirs for our white light curve fit. For consistency with Madhusudhan et al. (2023), we experiment with including both a linear and quadratic term in time, but we find that only a linear term is favored; we therefore choose to only include a linear term for our systematics model. We show the results of our white light curve fit in Figure 2.
@@
For our spectrophotometric fit, we fix the limb darkening parameters and linear systematic term to the fitted white light curve values in addition to the already fixed orbital parameters. We elect to fix the spectrophotometric limb darkening values to the white light curve values as limb darkening is not expected to be strongly wavelength dependent at the redder wavelengths of NIRSpec (May & MacDonald et al. 2023 2023; Moran & Stevenson et al. 2023 2023). We account for red noise in our spectrophotometric fit by inflating error bars in quadrature for wavelength bins where the residual RMS-bin size trend is above log linear. For most points this does not result in error bar inflation, and for points that are inflated, it is typically by a factor of no more than two. To demonstrate the fidelity of our fitting procedure given that FIREFLy is proprietary software, we show all of the spectroscopic light curve fits and residuals for each bin of our reduction in Appendix B.
@@
To reduce the NIRSpec/G395H observations, we also use the exoTEDRF pipeline (Radica et al., 2023; Feinstein et al., 2023; Radica, 2024), which has recently been updated to support NIRSpec observations (Radica, 2024, Ahrer et al. submitted). We start from the raw, uncalibrated data files and follow the procedure outlined in Ahrer et al. (submitted). For completeness, we summarize the key points here. We perform standard saturation and superbias corrections on the data frames (e.g., Alderson et al., 2023) before correcting the background and 1/ noise at the group-level, using the median method and a trace mask width of 16 pixels. As with SOSS, we perform a time-domain outlier flagging (Radica et al., 2024b) with a rejection threshold of 12. We then repeat the background and 1/ correction at the integration-level, after ramp fitting, in order to remove any remaining traces of the background flux. We then interpolate all pixels which have a non-zero data quality flag, or are flagged as 10 temporal outliers using a median of the surrounding pixels in space and time, respectively. For the spectral extraction, we use the edgetrigger algorithm (Radica et al., 2022b) to determine the centroids of the NRS1 and NRS2 spectral traces, and then perform a box aperture extraction with a width of 8 pixels.
We follow the same light curve fitting procedure as with the SOSS light curves — jointly fitting the NRS1 and NRS2 white light curves, sharing the planet’s orbital parameters between the two and independently fitting for chromatic parameters like the scaled planet radius, limb darkening and systematics. As with SOSS, we fix the orbital parameters to the values used by Madhusudhan et al. (2023) (there is, once again, no difference in the resulting spectra using the Radica et al. (2022a) orbital solution). There is no spot crossing in the NIRSpec light curves, and so we do not include a Gaussian spot model. However, we keep the linear slope and error inflation term added in quadrature to the flux errors that were included in the SOSS fits. In all, our white light curve fits have 15 free parameters. We again use emcee for the fits, using the same amount of chains, steps, and burn in as above.
@@ 3.3 Eureka! Reductions
For our first Eureka! reduction, which we henceforth refer to as reduction “A”, we start with the uncalibrated files and run Stage 1 and Stage 2 which are wrapped around the jwst pipeline. We use the default steps in both stages with the exception of the jump step and the photom step, which we skip. We also use Eureka!’s group-level background subtraction and a custom bias scale factor (using a smoothing filter with a window length of 30 integrations and group 1) in Stage 1444See also Moran & Stevenson et al. 2023 2023.
@@
In Stage 3 we extract the stellar time-series spectra. First, we correct the curved trace of the spectra and perform a background subtraction where we used the median of the entire frame, excluding the area within 5 pixels of the center of the trace. Prior to the calculation of the median background, we perform two iterations of outlier masking using thresholds of 3 along the time and 5 times the median along the spatial axis. We use an aperture of 7 pixels, corresponding to a 3-pixel half width, to extract the time-series spectra using the optimal extraction method (Horne, 1986).
@@
We generate pixel-level and binned light curves in Stage 4 of Eureka! following the same binning scheme as FIREFLy. We use a 3-clipping compared to a 10-pixel-rolling-median to reduce the number of outliers. We fit the light curves by fixing the system parameters to those reported in Madhusudhan et al. (2023) to allow for direct comparison, and we used the MCMC python package emcee. We opt to use a quadratic limb-darkening law, where we fix one parameter () to limb darkening parameters generated using the exotic-ld Python package (Grant & Wakeford, 2024) and the MPS-ATLAS (set 2) stellar atmosphere model grid (Kostogryz et al., 2023). We use a simple linear trend in time to account for a systematic trend in our light curves.
@@
Our second Eureka! reduction, which we will henceforth refer to as reduction “B”, used Eureka! version 0.11.dev446+gf5d684ee.d20240712; we started with the _uncal files produced with SDP_VER 2023_3b, and we used version 1.15.1 of the jwst pipeline with CRDS version 11.17.26 and CRDS context jwst_1252. Our reduction procedure largely followed the Eureka! reduction described in Schlawin et al. (2024); we summarize the important differences between this Eureka! reduction and that of Schlawin et al. (2024) below.
@@
We increased the Stage 1 jump step rejection threshold to 15 (instead of 6). In Stage 3, we 4-clipped background pixels along the time axis (instead of 10) and subtracted the mean background flux per column and per integration computed using only the pixels 6 px away from the center of the spectral trace (instead of 7 px). Our optimal spectral extraction (Horne, 1986) used a spatial profile with a half-width of 4 px (instead of 5 px). In Stage 4, we manually masked several pixel columns (18 in NRS1 and 9 in NRS2) that exhibited unusually high noise levels compared to their neighbors, likely due to unmasked bad pixels in an earlier stage. We then spectrally binned the data into multiple different spectral resolutions and masked any integration that was a 4 outlier (instead of 3.5) with respect to a 20-integration wide boxcar filter, likely due to a missed cosmic ray.
@@
When fitting the light curves, our astrophysical model consisted of a batman transit model with quadratic limb-darkening coefficients fixed to a PHOENIX limb-darkening model computed by exotic-ld (Husser et al., 2013; Grant & Wakeford, 2024). We first verified that our broadband light curves were consistent with the orbital parameters reported by Madhusudhan et al. (2023), with our joint fit of the NRS1 and NRS2 broadband lightcurves finding BMJD_TDB, , and (we had fixed the orbital period to that of Benneke et al. 2019b). We then fixed our orbital parameters to those of Madhusudhan et al. (2023). Our systematic model consisted of a linear trend in time and a linear correlation with the position and the width of the spatial PSF. We also fitted for a white noise multiplier to account for any excess white noise in the data beyond the estimated photon noise. We trimmed the first 25 integrations to remove any initial detector settling. To estimate the best-fit model values and their corresponding uncertainties, we used the dynesty nested sampling algorithm (Speagle, 2020) with 121 live points, ‘multi’ bounds, the ‘rwalk’ sampling algorithm, and a convergence criterion of , where is the Bayesian evidence.
@@ 4.1 Retrieval Configuration
We employ two open source retrieval codes, POSEIDON and BeAR, to explore the range of atmospheric properties consistent with the revised K2-18 b’s transmission spectra derived above. We additionally show retrieval results using the data from Madhusudhan et al. (2023) in Appendix A. Our retrieval configurations are constructed to encompass the parameter space explored by Madhusudhan et al. (2023), while also considering additional factors (such as planetary mass uncertainties and including the 2nd order NIRISS SOSS data). Our model settings, retrieval configurations555In this case, the “model” describes the structure of the planet’s atmosphere in the context of an observed transmission spectrum, with different potential “assumptions/settings” made for it. These include which potential aspects of the model to consider, the preferred parameterization of each component of the model, which parameters are fixed as well as the values they are fixed to, and the priors on each free parameter in the model. On the other hand, the “retrieval framework” describes how the model is fit to the data. This includes the way that the planet’s atmosphere and transmission through it are numerically modeled, the data used in these numerical models, and the code used to explore the parameter space. The combination of these is what we describe as the overall “retrieval configuration.”, and a comparison with Madhusudhan et al. (2023) are summarized in Table 3.
@@
| (citation) | Benneke et al. (2019a) | Benneke et al. (2017) | Benneke et al. (2019a) |
@@ 4.1.1 POSEIDON
We first retrieved K2-18 b’s transmission spectrum using the retrieval code POSEIDON666https://github.com/MartianColonist/POSEIDON (MacDonald & Madhusudhan, 2017; MacDonald, 2023). POSEIDON is widely applied to interpret JWST transmission spectra, including hot Jupiters (e.g., Grant et al., 2023; Fournier-Tondreau et al., 2024b), sub-Neptunes (Piaulet-Ghorayeb et al., 2024), super-Earths (e.g., Moran & Stevenson et al. 2023 2023; May & MacDonald et al. 2023 2023), and terrestrial worlds (e.g., Lim et al., 2023; Cadieux et al., 2024). Here we use POSEIDON to conduct a systematic retrieval survey of the sensitivity of K2-18 b’s atmospheric inferences to 60 different data variants (reduction codes and/or data spectral resolution). Our POSEIDON retrieval configuration is constructed to be as close as possible to that described in Madhusudhan et al. (2023), allowing a fair comparison, but we emphasize and justify below several areas where we adopted different model choices.
@@
Our atmosphere model for K2-18 b assumes a H2-He background gas (with a solar ratio of He/H2 = 0.17; Asplund et al. 2009), commensurate with its planetary radius and mass ( and , respectively). We construct atmospheres ranging from –10 bar with 100 layers spaced uniformly in log-pressure under hydrostatic equilibrium. We adopt bar as the lowest atmosphere pressure (vs. the bar limit implied by Madhusudhan et al. 2023) to reduce the risk of strong bands saturating the atmosphere. We solve hydrostatic equilibrium using a boundary condition where the freely retrieved reference pressure corresponds to the radial distance where (the radius from Benneke et al. 2019a). The atmospheric pressure-temperature (P-T) profile is freely retrieved following the prescription in Madhusudhan & Seager (2009). We consider the same potential molecular species as Madhusudhan et al. (2023), which includes the major O-, C-, and N-bearing species expected at K2-18 b’s equilibrium temperature (255 K; Benneke et al. 2019a): H2O, CH4, NH3, HCN, CO, and CO2; as well as several disequilibrium species proposed as biosignatures in cool sub-Neptunes (Madhusudhan et al., 2021): C2H6S (DMS), CS2, CH3Cl, OCS, and N2O. We fit the volume mixing ratios of these gases independently, with their number densities then determined via the ideal gas law. Given the large uncertainties on K2-18 b’s mass (; Cloutier et al. 2019), we prescribe the planet’s mass as a free parameter using a Gaussian prior (Madhusudhan et al. 2023 fixed K2-18 b’s mass), allowing mass uncertainties to propagate into the retrieved atmospheric composition (Batalha et al., 2019a). Finally, we consider aerosols following the inhomogeneous cloud and haze parameterization of MacDonald & Madhusudhan (2017).
@@
We calculate model spectra by solving the equation of radiative transfer in a cylindrical coordinate system for 100 incident stellar rays attenuated by atmospheric opacity along the line of sight. The POSEIDON opacity database recently underwent a major update in POSEIDON v1.2 (Mullens et al., 2024), including a state-of-the-art new ExoMol CH4 line list (Yurchenko et al., 2024). Our POSEIDON opacities are pre-computed at a high resolution ( = 0.01 cm-1, equivalent to at 1 ) for a grid of temperatures and pressures using the open source Python package Cthulhu777https://github.com/MartianColonist/Cthulhu (Agrawal & MacDonald, 2024). Our K2-18 b retrievals use opacities derived from the following line lists/measured cross sections: H2O (POKAZATEL; Polyansky et al. 2018), CH4 (MM; Yurchenko et al. 2024), NH3 (CoYuTe; Coles et al. 2019), HCN (Harris; Barber et al. 2014, CO (Li2015; Li et al. 2015), CO2 (UCL-4000; Yurchenko et al. 2020), C2H6S (DMS) (HITRAN measured xsec; Sharpe et al. 2004) CS2 (HITRAN-2020; Gordon et al. 2022), CH3Cl (HITRAN-2020; Gordon et al. 2022), OCS (OYT8; Owens et al. 2024), and N2O (HITEMP; Hargreaves et al. 2019). We further include continuum collision-induced absorption from H2-H2 and H2-He pairs (Karman et al., 2019) and Rayleigh scattering for all gases (MacDonald & Lewis, 2022). We calculate model transmission spectra using opacity sampling on a wavelength grid from 0.58–5.3 at a spectral resolution of 100,000. We adopt this exceptionally high spectral resolution for all our POSEIDON retrievals for three reasons: (i) to ensure negligible opacity sampling errors when interpreting pixel-level data; (ii) to consistently sample cross sections for every data combination (even at lower resolutions); and (iii) to reduce the risk of opacity accuracy differences between our work and Madhusudhan et al. (2023) (since the latter did not state their assumed sampling resolution). While our 100,000 models impose a greater computational burden than is strictly necessary for retrievals of JWST data888Through the JWST Transiting Exoplanet Early Release Science Program, we found a useful heuristic that 20,000 suffices for = 100 and 60,000 suffices for 3,000 (pixel-level)., this choice ensures high accuracy in our retrieved atmospheric properties and reduces the risk of an insufficiently high opacity resolution inconsistent with the methodology of Madhusudhan et al. (2023).
@@
Our POSEIDON retrievals additionally account for stellar contamination from unocculted active regions, commonly known as the transit light source effect (Rackham et al., 2018). We employ a one-heterogeneity model (e.g., Rathcke et al., 2021; Lim et al., 2023; Fournier-Tondreau et al., 2024a), defined by the stellar photosphere temperature, the heterogeneity temperature (corresponding to spots/faculae if the heterogeneity is cooler/warmer than the photosphere), and the heterogeneity covering fraction. The stellar contamination contribution is calculated by interpolating PHOENIX models (Husser et al., 2013) using the PyMSG package (Townsend & Lopez, 2023). For fixed stellar properties, we adopt the stellar radius from Benneke et al. (2019a) () and the stellar surface gravity, metallicity, and effective temperature from Tabernero et al. (2024) ([Fe/H] = , = 4.93, = 3563 K). These updated K2-18 stellar properties were published after Madhusudhan et al. (2023), so we adopt them for a more accurate characterization of the transit light source effect.